OpenStudy (anonymous):

Need help determining the missing steps in my professor's notes. Coded steps soon to come

OpenStudy (anonymous):

\[\large{\tau_{rz}={-\mu \over g_c}({du \over dr})_{r=r_i}}\]where mu, gc, and ri are constants\[u=u_{\max}[1-({r \over r_i})^2]\]\[\large\tau_{rz}={-\mu \over g_c} \times {u_\max \over r_i} \times -2r; \text{for } {r=r_i}\]\[\large \tau_{rz}={2\mu u_\max \over g_c r_i}\]

OpenStudy (anonymous):

i'm confused after he substitutes 'u' into equation 1 to achieve the new Tau equation in line 3

OpenStudy (anonymous):

I'm guessing he takes the derivative of u then substitutes in for du

OpenStudy (anonymous):

u(max) is also a constant

ganeshie8 (ganeshie8):

\(\large \mathbb{u=u_{\max}[1-({r \over r_i})^2]}\)

ganeshie8 (ganeshie8):

differentiate both sides with.respect.to r

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{du}{dr}= \frac{d}{dr} (u_{\max}[1-({r \over r_i})^2])}\)

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{du}{dr}= \frac{d}{dr} (u_{\max}[1-({r \over r_i})^2])}\) \(\large \mathbb{\frac{du}{dr}= \frac{d}{dr} (u_{\max}-u_{\max}({r \over r_i})^2)}\)

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{du}{dr}= \frac{d}{dr} (u_{\max}[1-({r \over r_i})^2])}\) \(\large \mathbb{\frac{du}{dr}= \frac{d}{dr} (u_{\max}-u_{\max}({r \over r_i})^2))}\) \(\large \mathbb{\frac{du}{dr}= \frac{d}{dr} (u_{\max}-\frac{u_{\max}}{r_i^2}(r^2)})\)

ganeshie8 (ganeshie8):

derivative of 1 is 0

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{du}{dr}= \frac{d}{dr} (u_{\max}[1-({r \over r_i})^2])}\) \(\large \mathbb{\frac{du}{dr}= \frac{d}{dr} (u_{\max}-u_{\max}({r \over r_i})^2))}\) \(\large \mathbb{\frac{du}{dr}= \frac{d}{dr} (u_{\max}-\frac{u_{\max}}{r_i^2}(r^2)})\) \(\large \mathbb{\frac{du}{dr}= 0-\frac{u_{\max}}{r_i^2}(2r)}\) \(\large \mathbb{\frac{du}{dr}= \frac{u_{\max}}{r_i^2}(2r)}\)

ganeshie8 (ganeshie8):

^^you should get that

OpenStudy (anonymous):

\[\large du={2u_{\max}r \over r_i^2}\]

ganeshie8 (ganeshie8):

dont ignore the dr in the bottom of du ok ?

ganeshie8 (ganeshie8):

its du/dr not du

OpenStudy (anonymous):

Can you help me explain the terminolgy a little bit. You are leaving it as du/dr because it is with respect to r

ganeshie8 (ganeshie8):

both are different things one is called derivative, the other is called differential

OpenStudy (anonymous):

d/dr just means the derivative with respect to r as du/dr means the derivative of u with respect to r

ganeshie8 (ganeshie8):

yup !

OpenStudy (anonymous):

du/dr would be the differential

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{du}{dr}= \frac{d}{dr} (u_{\max}[1-({r \over r_i})^2])}\) \(\large \mathbb{\frac{du}{dr}= \frac{d}{dr} (u_{\max}-u_{\max}({r \over r_i})^2))}\) \(\large \mathbb{\frac{du}{dr}= \frac{d}{dr} (u_{\max}-\frac{u_{\max}}{r_i^2}(r^2)})\) \(\large \mathbb{\frac{du}{dr}= 0-\frac{u_{\max}}{r_i^2}(2r)}\) \(\large \mathbb{\frac{du}{dr}= \frac{u_{\max}}{r_i^2}(2r)}\) \(\large \mathbb{\frac{du}{dr}\Big|_{r=r_i} = \frac{u_{\max}}{r_i^2}(2r_i) = \frac{u_{\max}}{r_i}(2) }\)

OpenStudy (anonymous):

so now the professor probably moved the dr to the other side and substiuted du then reduced to the final equeation

ganeshie8 (ganeshie8):

differentials : \(du, dr\) derivatives : \(\frac{du}{dr}, \frac{d}{dr}\)

ganeshie8 (ganeshie8):

ur professor simply substituted the value of \(\frac{du}{dr}\)

OpenStudy (anonymous):

Thanks for your help. Makes sense now.

ganeshie8 (ganeshie8):

good, u wlc =)